# Properties and dynamics of generalized squeezed states

### v1

@@ V Conclusion
We have demonstrated, using a variety of numerical simulations and additional numerical checks, that generalized squeezing with $n\geq 3$ is qualitatively different from the lower-order operations of displacement and two-photon squeezing. The indefinite growth for the low-order operations is replaced with oscillatory dynamics for higher-order operations. Although some quantities exhibit finite-size effects in our simulations, we believe that several features in the results point to the oscillations being a physical phenomenon that should be observable in future experiments. As experimental capabilities on nonlinear quantum optics are advancing rapidly, we expect that our results will serve as an important reference point for the development of generalized squeezing protocols in the future.
@@ Acknowledgment
This work was supported by Japan’s Ministry of Education, Culture, Sports, Science and Technology (MEXT) Quantum Leap Flagship Program Grant Number JPMXS0120319794.

### v2

@@ V Conclusion
We have demonstrated, using a variety of numerical simulations and additional numerical checks, that generalized squeezing with $n\geq 3$ is qualitatively different from the lower-order operations of displacement and two-photon squeezing. The indefinite growth of the photon number for the low-order operations is replaced with oscillatory dynamics for higher-order operations. Although some quantities exhibit finite-size effects and would extrapolate to infinity for an infinite simulation size, several features in the results point to the oscillations being a physical phenomenon that is mathematically well behaved and should be observable in future experiments. As experimental capabilities on nonlinear quantum optics are advancing rapidly, we expect that our results will serve as an important reference point for the development of generalized squeezing protocols in the future.
@@ Acknowledgment
We would like to thank Daniel Braak, Suguru Endo, Lin Tian and Fumiki Yoshihara for useful discussions. This work was supported by Japan’s Ministry of Education, Culture, Sports, Science and Technology (MEXT) Quantum Leap Flagship Program Grant Number JPMXS0120319794.
Data availability
The datasets generated and/or analysed during the current study are available from the corresponding author on reasonable request.